OpenStudy (anonymous):

the ratio of girls to boys at school is 7:4. if there are currently 189 girls at the school, then how many additional boys would be needed so that the resulting ratio of girls to boys is 3:2? Show answer in 2 different ways... please explain.

OpenStudy (anonymous):

So if the number of boys = x 7/4 x =189 Solve for the initial number of boys x= 189 / (7/4) x=108 Then how many more would you need for a 3/2 ratio? hint: 3/2=189/ new number of boys

OpenStudy (anonymous):

is there another way of solving?

OpenStudy (anonymous):

is there another way of solving?

OpenStudy (anonymous):

You could do things in a different order, or maybe try to make an equation to solve all of it at once which would just be more complicated..